OpenStudy (cherie_magee):

Write the following equation in the general form Ax + By + C = 0. y - x - 1 = 0 -2x + 3y - 6 = 0 2x - 3y + 6 = 0 2x - 3y - 6 = 0

OpenStudy (primeralph):

Is it y - x - 1 = 0 or y/3 - x/2 - 1 = 0 ?

OpenStudy (cherie_magee):

im not sure that's what i'm trying to figure out @primeralph

OpenStudy (primeralph):

I'm asking if that's the question. Your question seems wrong.

OpenStudy (cherie_magee):

i dont know thats just how they have it on here i'm confused af @primeralph

OpenStudy (primeralph):

Can you post a screenshot of the question?

OpenStudy (cherie_magee):

Write the following equation in the general form Ax + By + C = 0. 1/2y-1/3x-1=0 -2x + 3y - 6 = 0 2x - 3y + 6 = 0 2x - 3y - 6 = 0

OpenStudy (cherie_magee):

@primeralph

OpenStudy (primeralph):

That's much better. C.

OpenStudy (primeralph):

Just multiply it by 6.